OpenStudy (anonymous):

4^x-3*2^(x+2)+2^5=0 solve for x

OpenStudy (anonymous):

(4^x/4^3)+(2^x,2^2)+2^5=0 (2^2x/64)+(4.2^x)+2^5=0 now u can take 2^x=u and it becomes a quadratic eq, which can be solved bt quadratic formula

OpenStudy (anonymous):